five years ago, Gemma was twice as old as tony is now, in seven years Gemma will be five times as old as Tony is now. What are Gemma and Tony's current age?
Answer by richwmiller(17219)

g-5=2t
g+7=5t
g=2t+5
g=5t-7
2t+5=5t-7
12=3t
4=t
g-5=8
g=13
check
13+7=5*4
20=20
ok